OpenStudy (anonymous):

Robert and Elena usually order a medium pizza for $14.00. The large pizza costs 40% more. What is the cost of the large pizza?

OpenStudy (anonymous):

\[$14\times.40=5.6\]

OpenStudy (anonymous):

\[$14+5.6=$19.6\]

OpenStudy (anonymous):

(14*40)/100=$5.6....... so the actual rate is=$14+5.6=19.6...